# On-device Sora: Enabling Diffusion-Based Text-to-Video Generation for Mobile Devices

### v1

@@ 8.4. Concurrent Inference with Dynamic Loading
Figure 14-(b) illustrates the model block loading and inference cycles of STDiT (Zheng et al., 2024) when applying the proposed Concurrent Inference (Sec. 6.1), whereas Figure 14-(a) depicts the case without its application. It can be observed that, with Concurrent Inference, the GPU executes each model block for inference without almost no idle time in parallel with the model block loading, indicated by the overlap between the red (loading) and black (inference) boxes in Figure 14-(b), resulting in a block inference latency reduction from 29s to 23s. In contrast, without Concurrent Inference, each model block inference is executed only after the corresponding block is fully loaded to memory, indicated by the lack of overlap between the red (loading) and black (inference) boxes in Figure 14-(a). Consequently, the total latency of the full denoising process using multiple executions of STDiT (Zheng et al., 2024) is reduced by approximately 25%, decreasing from 1,000 to 750 seconds for 30 denoising steps. Given that STDiT is executed multiple times to perform numerous denoising steps, it significantly accelerates the STDiT’s overall inference. In addition, when applied with Dynamic Loading (Sec. 6.2), it achieves an additional average speed improvement of 17 seconds as reloading is not required for certain model blocks that are retained in memory, provided in Equation 11.
Figure 15 shows the case of T5 (Raffel et al., 2020). Unlike STDiT, which exhibits similar latencies for both block loading and inference execution, T5 exhibits a much longer block loading latency relative to inference latency. Consequently, the overlap between the model loading and inference is minimal, as shown by the small region of overlap between the red (loading) and black (inference) boxes. As a result, the latency improvement is expected to be less substantial, as in Equation 10. Nevertheless, the inference latency is reduced from 164 to an average of 137 seconds, achieving a reduction of 16%. This result implies the effectiveness of concurrent loading and inference, even in cases when there is an imbalance between the latencies of model block loading and inference.
@@
Figure 14. The block loading and inference cycles of STDiT (Zheng et al., 2024) without (a) and with (b) Concurrent Inference. The red box represents the loading cycle, while the black box indicates the model block inference on the iPhone 15 Pro’s GPU.
@@
Figure 15. The block loading and inference cycles of T5 (Raffel et al., 2020) without (a) and with (b) Concurrent Inference.
8.5. Video Generation Latency
Table 5 shows video generation latencies of two resolutions when each of the proposed methodologies—LPL, TDTM, and CI-DL—is applied individually, as well as the latency when all of them are applied together (‘All’). Latencies are measured with LPL activated at the 15th denoising step and TDTM applied throughout all steps, reported as the mean of three independent experiments. ‘STDiT’ and ‘Total’ specify whether the latency is measured solely for STDiT (Zheng et al., 2024) or for end-to-end video generation, including T5 (Raffel et al., 2020) and VAE (Doersch, 2016). The results demonstrate substantial latency reductions for each methodology compared to the case without using the proposed methods, i.e., 293.51 vs. 1768.32 seconds (Table 1) for STDiT. For the 192$\times$192 resolution video, STDiT (denoising process) takes less than five minutes when all three methodologies are applied. Additionally, it indicates that the methodologies do not interfere with each other, instead work synergistically to enhance latency.
Table 5. Ablation study on video generation latency (s). ‘All’ denotes the combined application of LPL, TDTM, and CI-DL.
Resolution Measurement LPL TDTM CI-DL All
192$\times$192 STDiT 390.72 696.03 566.81 293.51
Total 514.24 823.03 691.25 416.50
256$\times$256 STDiT 573.88 965.40 947.67 454.48
Total 754.08 1148.63 1127.88 638.09
9. Discussions and Limitations
@@
Latency Improvement. Although On-device Sora enables efficient video generation, the generation latency remains higher compared to utilizing high-end GPUs; it requires several minutes to generate a video, whereas an NVIDIA A6000 GPU takes one minute. This discrepancy is evident due to the substantial disparity in computational resources between them. For instance, the iPhone 15 Pro’s GPU features up to 2.15 TFLOPS with 3.3 GB of available memory, compared to the NVIDIA A6000, which offers up to 309 TFLOPS and 48 GB of memory. Despite this significant resource gap, On-device Sora achieves exceptional efficiency in video generation. Currently, it utilizes only the iPhone 15 Pro’s GPU. We anticipate that the latency could be significantly enhanced if it can leverage NPU (Neural Processing Unit), e.g., the iPhone 15 Pro’s Neural Engine (Apple, 2023), which delivers a peak performance of 35 TOPS. However, the current limitations in Apple’s software and SDK support for state-of-the-art diffusion-based models (Zheng et al., 2024) make the iPhone’s NPU challenging to utilize effectively. We look forward to the development of software support for NPUs and leave the exploration of this for future work. Also, we plan to investigate the potential of NPUs on a variety of mobile devices, such as Android smartphones.
@@
Model Optimization. In On-device Sora, only T5 (Raffel et al., 2020) is quantized to int8, reducing its size from 18 GB to 4.64 GB, while STDiT (Zheng et al., 2024) and VAE (Doersch, 2016) are executed with float32 due to their performance susceptibility, which has a significant impact on video quality. Additionally, we do not apply pruning (Reed, 1993) or knowledge distillation (Gou et al., 2021), as these methods also drop visual fidelity. In particular, we observe that naively shrinking STDiT (Zheng et al., 2024) leads to significant visual loss, caused by iterative denoising steps, where errors propagate and accumulate to the final video. Another practical difficulty in achieving lightweight model optimization is the lack of resources required for model optimization; both re-training and fine-tuning state-of-the-art diffusion-based models typically demand several tens of GPUs, and the available datasets are often limited to effectively apply optimization methods. To tackle this challenge, we propose model training-free acceleration techniques for video generation in this work, i.e., Linear Proportional Leap (Sec. 4) and Temporal Dimension Token Merging (Sec. 5). We recognize further model optimization as a potential direction for our future work.
Image-to-Video Generation. Building on the proposed On-device Sora, it is also feasible to extend its capabilities to image-to-video generation, empowering users to create personalized videos based on their own visual data. Furthermore, On-device Sora could evolve to accept both text and images as a multi-modal generative solution, providing an integrated approach to versatile data generation, expected to facilitate the personalization of video generation on mobile devices. We envision that On-device Sora would lay the foundation for future multi-modal visual and textual generation applications on various mobile systems, fostering the on-device revolution and expansion of generative technologies.
10. Related Work
On-device Video Generation. Numerous video diffusion models have advanced the related technologies, including Make-A-Video (Singer et al., 2022), Snap Video (Menapace et al., 2024), Imagen Video (Ho et al., 2022a), Video Diffusion (Ho et al., 2022c), Stable Video Diffusion(Blattmann et al., 2023a), Dreamix (Molad et al., 2023), MCVD (Voleti et al., 2022), Tune-A-Video (Wu et al., 2023), Text2Video-Zero (Khachatryan et al., 2023), among others. Given the substantial resource requirements of video generation, some studies have investigated optimization under various conditions. For example, QVD (Tian et al., 2024) employs post-training quantization (Liu et al., 2021b) to enhance computational efficiency and reduce memory usage, while MagicVideo (Zhou et al., 2022) minimizes computations by generating videos in the latent space. CMD (Yu et al., 2024) decomposes videos into content frames and low-dimensional motion representations to reduce the computational cost of video generation. Similarly, SimDA (Xing et al., 2024) improves the efficiency of diffusion models by utilizing an effective adapter mechanism. While those studies have sought to enhance the efficiency of video generation, none of them have successfully demonstrated to generate videos on commodity mobile devices, primarily due to high resource demands even after applying their optimization techniques. To the best of our knowledge, the proposed On-device Sora is the first standalone and efficient video generation solution capable of operating on resource-constrained smartphones.
On-device Image Generation. There are many works that generate images on the device (Li et al., 2024b; Vasu et al., 2023; Castells et al., 2024; Jia et al., 2023; Belousov, 2021). However, on-device video generation is struggling due to various challenges. First, video generation requires processing substantially more complex and larger data than image generation (Blattmann et al., 2023a; Chen et al., 2017; Nan et al., 2024). While image generation focuses on a single frame, video generation involves handling multiple frames, with the data requirements scaling with the video’s frame rate (fps) and duration. Additionally, video generation must maintain temporal consistency across frames (Menapace et al., 2024; Singer et al., 2022; Xing et al., 2024), i.e., it demands maintaining coherent temporal relationships, including object movement, lighting transitions, and background dynamics between successive frames. Achieving this level of consistency necessitates significantly more computation and memory. Furthermore, these requirements often require complex and large models, exacerbating the challenges when operating on resource-limited mobile devices.
Rectified Flow. While Open-Sora (Zheng et al., 2024) reduces the number of denoising steps by leveraging Rectified Flow (Liu et al., 2022), most related approaches (Esser et al., 2024; Zhu et al., 2025) require conditioned model training and/or distillation (Zhu et al., 2025). In contrast, the proposed Linear Proportional Leap effectively reduces the denoising steps without a significant performance drop, as validated using VBench (Huang et al., 2024), without requiring model re-training or distillation. Notably, it can be easily activated at runtime by just calculating the cosine similarities of drifts between consecutive denoising steps. By developing the current Euler method (Chen et al., 2018) into more advanced methodologies, it may be possible to further reduce denoising steps while ensuring stability within boundaries. We leave this for future work.
Token Merging. Most token merging methods (Bolya et al., 2022, 2023; Feng and Zhang, 2023) primarily focus on image generation, where tokens are merged based on the spatial similarity rather than temporal similarity. Although some temporal token merging techniques have been proposed, they are applied to models in other domains (Götz et al., 2024; Li et al., 2024a), not in video generation. As such, Temporal Dimension Token Merging is the first to apply token merging based on the successive similarities between frames in video generation. Additionally, while previous works predominantly apply token merging to self-attention due to performance degradation (Bolya et al., 2022, 2023; Feng and Zhang, 2023; Li et al., 2024a), On-device Sora shows that token merging can be applied to cross-attention with minimal performance loss, achieving 50% merging ratio.
11. Conclusion
We propose On-device Sora, the first pioneering solution for generating videos on mobile devices using diffusion-based models, which addresses several key challenges in video generation to enable efficient on-device operation. The issue of extensive denoising steps is addressed through Linear Professional Leap, the challenge of handling large tokens is mitigated with Temporal Dimension Token Merging, and memory limitations are overcome through Concurrent Inference with Dynamic Loading. These proposed efficient on-device video generation methodologies are not limited to On-device Sora but are broadly applicable to various applications, providing significant advancements in on-device video generation without additional model re-training.

### v2

@@ B.2 Concurrent Inference with Dynamic Loading
Fig. 12-(b) illustrates the model block loading and inference cycles of STDiT [83] when applying the proposed Concurrent Inference (Sec. A.1), whereas Fig. 12-(a) depicts the case without its application. It can be observed that, with Concurrent Inference, the GPU executes each model block for inference without almost no idle time in parallel with the model block loading, indicated by the overlap between the red (loading) and black (inference) boxes in Fig. 12-(b), resulting in a block inference latency reduction from 29s to 23s. In contrast, without Concurrent Inference, each model block inference is executed only after the corresponding block is fully loaded to memory, indicated by the lack of overlap between the red (loading) and black (inference) boxes in Fig. 12-(a). Consequently, the total latency of the full denoising process using multiple executions of STDiT [83] is reduced by approximately 25%, decreasing from 1,000 to 750 seconds for 30 denoising steps. Given that STDiT is executed multiple times to perform numerous denoising steps, it significantly accelerates the STDiT’s overall inference. In addition, when applied with Dynamic Loading (Sec. A.1), it achieves an additional average speed improvement of 17 seconds as reloading is not required for certain model blocks that are retained in memory, provided in Eq. 11.
@@
Figure 12: The block loading and inference cycles of STDiT [83] without (a) and with (b) Concurrent Inference. The red box represents the loading cycle, while the black box indicates the model block inference on the iPhone 15 Pro’s GPU.
@@
Figure 13: The block loading and inference cycles of T5 [48] without (a) and with (b) Concurrent Inference.
Fig. 13 shows the case of T5 [48]. Unlike STDiT, which exhibits similar latencies for both block loading and inference execution, T5 exhibits a much longer block loading latency relative to inference latency. Consequently, the overlap between the model loading and inference is minimal, as shown by the small region of overlap between the red (loading) and black (inference) boxes. As a result, the latency improvement is expected to be less substantial, as in Eq. 10. Nevertheless, the inference latency is reduced from 164 to an average of 137 seconds, achieving a reduction of 16%. This result implies the effectiveness of concurrent loading and inference, even in cases when there is an imbalance between the latencies of model block loading and inference.
B.3 Ablation Study of Linear Proportional Leap (LPL) on Other Text-to-Video Models
The primary constraint for applying Linear Proportional Leap (LPL) is that the model must be a flow-matching–based model trained with a rectified flow target [38]. Since Pyramidal Flow [29], one of the state-of-the-art open-source video generation models, satisfies this condition, we conduct experiments on it.
In a manner similar to Open-Sora with Rectified Flow [38], Pyramidal Flow consistently demonstrates a high degree of straightness in its generation process following the initial generation unit (as illustrated in Fig. 14). As demonstrated in Tab. 7, the application of LPL during generation reduces the overall generation time by nearly 38%, decreasing from 264.60 seconds to 165.15 seconds. Notably, Pyramidal Flow employs a cascading model procedure in which subtle variances are treated as noise to be subsequently denoised. Since these variances become negligible in the final output, Linear Proportional Leap is particularly well-suited to this architecture. Our experiments involving multiple prompts (as shown in Fig. 15) confirmed that removing nearly half of the total denoising steps does not compromise performance while preserving both visual quality and temporal consistency. Furthermore, we observed that additional steps at later units can also be omitted due to the high straightness of Pyramidal Flow. Future work will further investigate this phenomenon by exploring the compatibility of flow-based models with the current LPL algorithm, developing the algorithm to be more compatible to other type of models.
Models Time (s) Resolution DiT Steps Speedup
Pyramidal Flow - 2B 262.03 1280 $\times$ 768 270 1.00$\times$
Pyramidal Flow - 2B - LPL (2) 165.15 1280 $\times$ 768 159 1.59$\times$
Pyramidal Flow - 2B 48.88 640 $\times$ 384 270 1.00$\times$
Pyramidal Flow - 2B - LPL (2) 30.54 640 $\times$ 384 159 1.60$\times$
Table 7: Ablation study conducted on Pyramidal Flow [29] based on the miniFLUX [32] architecture. ‘DiT steps’ represent the total number of DiT forward computations required for the denoising process. Speedups are measured relative to the corresponding model, with execution time computed as the average of three independent runs. Notably, LPL does not degrade video quality, even when using nearly half of its total steps.
Figure 14: Cosine similarities between adjacent drifts estimated by Pyramidal Flow [29] at each stage. (a) and (b) show cosine similarities computed at the first and last units, respectively. The values approach 1.0 at the last unit across all stages. Computations follow the experimental settings in [29], using 20 denoising steps in (a) and 10 in (b).
Prompt: ”A movie trailer featuring the adventures of a 30-year-old spaceman wearing a red wool-knitted motorcycle helmet, blue sky, salt desert, cinematic style, shot on 35mm film, vivid colors.”
Prompt: ”A cybernetic humanoid scans the streets with red eyes as holographic screens flicker around its head, displaying futuristic data.”
Prompt: ”A green-eyed woman with auburn hair and a mysterious smile, wearing a flowing red dress.”
Figure 15: Videos generated using naive Pyramidal Flow [29] (top) and Pyramidal Flow with the proposed Linear Proportional Leap (bottom) under identical prompts. Linear Proportional Leap (LPL) reduces the total number of denoising steps by nearly half (i.e., from 270 to 159), while maintaining video generation results that are highly comparable to those achieved using the full set of denoising steps and preserving most details.
Appendix C Appendix - Discussions and Limitations
@@
Latency Improvement. Although On-device Sora enables efficient video generation, the generation latency remains higher compared to utilizing high-end GPUs; it requires several minutes to generate a video, whereas an NVIDIA A6000 GPU takes one minute. This discrepancy is evident due to the substantial disparity in computational resources between them. For instance, the iPhone 15 Pro’s GPU features up to 2.15 TFLOPS with 3.3 GB of available memory, compared to the NVIDIA A6000, which offers up to 309 TFLOPS and 48 GB of memory. Despite this significant resource gap, On-device Sora achieves exceptional efficiency in video generation. Currently, it utilizes only the iPhone 15 Pro’s GPU. We anticipate that the latency could be significantly enhanced if it can leverage NPU (Neural Processing Unit), e.g., the iPhone 15 Pro’s Neural Engine [1], which delivers a peak performance of 35 TOPS. However, the current limitations in Apple’s software and SDK support for state-of-the-art diffusion-based models [83] make the iPhone’s NPU challenging to utilize effectively. We look forward to the development of software support for NPUs and leave the exploration of this for future work. Also, we plan to investigate the potential of NPUs on a variety of mobile devices, such as Android smartphones.
@@
Model Optimization. In On-device Sora, only T5 [48] is quantized to int8, reducing its size from 18 GB to 4.64 GB, while STDiT [83] and VAE [15] are executed with float32 due to their performance susceptibility, which has a significant impact on video quality. Additionally, we do not apply pruning [49] or knowledge distillation [19], as these methods also drop visual fidelity. In particular, we observe that naively shrinking STDiT [83] leads to significant visual loss, caused by iterative denoising steps, where errors propagate and accumulate to the final video. Another practical difficulty in achieving lightweight model optimization is the lack of resources required for model optimization; both re-training and fine-tuning state-of-the-art diffusion-based models typically demand several tens of GPUs, and the available datasets are often limited to effectively apply optimization methods. To tackle these challenges, we propose model training-free acceleration techniques for video generation in this work, i.e., Linear Proportional Leap (Sec. 4) and Temporal Dimension Token Merging (Sec. 5).
Straightness Constraints. For video generation models that do not exhibit straightness during their denoising procedures, such as CogVideoX [74], which employs DPM-Solver [82], Linear Proportional Leap (LPL) is currently inapplicable. In such cases, as described in previous work [75], an alternative approach involves predicting modifications to the model’s noise schedule by employing additional methodologies, such as reinforcement learning. However, these kind of methods require additional training which may invoke extra costs, and necessitates retraining whenever the target model architecture changes. By contrast, LPL is a plug-and-play algorithm that can be applied in a model-agnostic manner, as long as a sufficient level of straightness is maintained during the denoising process, thereby eliminating the need for additional training. In future work, we plan to investigate the optimal reduction point achievable with this algorithm, and we anticipate that it will evolve into a widely applicable methodology for all Rectified Flow-based models [38].